# Diversity of low-mass planet atmospheres in the C-H-O-N-S-Cl system with interior dissolution, non-ideality, and condensation: Application to TRAPPIST-1e and sub-Neptunes

### v2

@@ Appendix E Treatment of oxygen
The condensed reservoir (molten/solid rock) of a rocky planet dominates its gaseous reservoir in terms of mass or number of moles. During the magma ocean stage, in which the molten mantle and atmosphere communicate due to similar dynamic timescales, it is therefore expected that mantle chemistry plays a role in regulating the chemistry of the combined atmosphere-mantle system. In particular, oxygen fugacity (O2) is a crucial parameter in mantle geochemical systems because it governs the oxidation state of elements, influencing mineral stability, phase equilibria, and the speciation of volatiles such as hydrogen, carbon, and sulfur. It is reasonable to assume that oxygen fugacity (O2) is buffered by the iron-wüstite (IW) equilibrium in the early history of a rocky planet because planetary accretion and differentiation involve metal-silicate equilibration under reducing conditions.
Rock analyses confirm the redox state of Earth’s magma ocean was close to the IW buffer (Sossi et al., 2020). Hence for the high temperature calculations of both TRAPPIST-1e and K2-18b the control of mantle chemistry on atmospheric speciation is encapsulated in the constraint on O2, which is defined relative to the IW buffer that exclusively determines (absolute) O2 (in bar) at a given temperature and a reference pressure of 1 bar. Representing O2 relative to the IW buffer does not necessitate that the atmosphere of K2-18b is actually buffered by the Fe–FeO reaction at present day. Furthermore, it should be appreciated that imposing a buffered oxygen constraint does not in itself constrain the total oxygen abundance. This is evident, for example, by considering the following redox reaction:
@@
where is temperature and fugacity, in which the later can be considered equivalent to partial pressure for this discussion. If O2 is constrained relative to the IW buffer then both O2 and are solely functions of temperature and these collectively govern the ratio of H2 to H2O but importantly not their absolute values. Imposing an additional hydrogen abundance constraint (or alternatively pressure constraint) is required to break the degeneracy to allow a unique determination of H2 and H2O. That is, 2 equations (Equation E3 and hydrogen mass balance) can be solved to determine 2 unknowns (H2 and H2O), implicitly constraining the total oxygen abundance. Hence if O2 is prescribed as a constraint, Atmodeller solves for the total oxygen abundance self-consistently as part of the solution process. For the cooled atmosphere calculations of TRAPPIST-1e, the elemental abundance of species in the atmosphere (including oxygen) are prescribed as constraints, in which case Atmodeller then ensures a self-consistent determination of O2.
Appendix F Interplay of solubility and non-ideality
Chachan & Stevenson (2018) and Kite et al. (2019) present models for H2 dissolution in sub-Neptunes, for ideal and non-ideal scenarios, respectively. Here, we develop a toy model to demonstrate how these previous results are also obtained with the formulation of Atmodeller. We start with a simplified mass balance equation used by Atmodeller (e.g., Appendix C, Bower et al., 2022), where the first term is the dissolved mass of volatile, the second term the mass of volatile in the atmosphere, and the third term the total mass of volatile:
@@ Appendix F Interplay of solubility and non-ideality
We employ a general expression for solubility with a power-law dependence on fugacity:
where , , and are constants specific to a particular equilibrium between a gas species and its dissolved component and is temperature at the atmosphere-interior interface. We define a functional form for the fugacity coefficient (Equation 1) that captures the general tendency of fugacity coefficients to increase exponentially with pressure:
where is total pressure, is a constant that regulates non-ideality and bar is the pressure at which the volatile behaves ideally. Hence ideal behaviour is recovered as or . For simplicity, temperature-dependence can also be accounted for through the choice of P0. Substituting Equations F2 and F3 into F1 gives:
where:
@@
The ratio of the terms in the square brackets corresponds to the ratio of the dissolved volatile mass and the atmospheric volatile mass :
where is the gravitational constant. For the limit of ideal behaviour () and a fully molten core () the scaling in Equation 12 of Chachan & Stevenson (2018) is recovered. Hence the dissolved volatile mass fraction relative to that of the atmosphere increases with planet size and higher surface temperature (Figure 1, Chachan & Stevenson, 2018; Gillmann et al., 2024). Kite et al. (2019) additionally include the non-ideal dissolution of H2 into magma driven by the non-linear increase of fugacity with pressure (quantified by ), which again gives rise to enhanced volatile storage in the interior.
@@ Appendix G Additional K2-18b Figures
Figures G1 and G2 show how atmospheric pressures (a), atmospheric C/H and C/O ratios by mass (b), solubilities (c) and fugacity coefficients (d) vary with oxygen fugacity and total C/H mass ratio, respectively, between the real and ideal cases. In these simulations, the planet mass, surface radius and temperature are the same as those in Section 3.2 and Figures 6 and 7, and only one input parameter is varied (either oxygen fugacity or C/H mass ratio) while the others are kept constant.

### v3

@@ Appendix E Treatment of oxygen
The condensed reservoir (molten/solid rock) of a rocky planet dominates its gaseous reservoir in terms of mass or number of moles. During the magma ocean stage, in which the molten mantle and atmosphere communicate due to similar dynamic timescales, it is therefore expected that mantle chemistry plays a role in regulating the chemistry of the combined atmosphere-mantle system. In particular, oxygen fugacity (O2) is a crucial parameter in mantle geochemical systems because it governs the oxidation state of elements, influencing mineral stability, phase equilibria, and the speciation of volatiles such as hydrogen, carbon, and sulfur. It is reasonable to assume that oxygen fugacity (O2) is buffered by the IW equilibrium in the early history of a rocky planet because planetary accretion and differentiation involve metal-silicate equilibration under reducing conditions.
Rock analyses confirm the redox state of Earth’s magma ocean was close to the IW buffer (Sossi et al., 2020). Hence, for the high-temperature calculations of both TRAPPIST-1e and K2-18b the control of mantle chemistry on atmospheric speciation is encapsulated in the constraint on O2, which is defined relative to the IW buffer that exclusively determines (absolute) O2 (in bar) at a given temperature and a reference pressure of 1 bar. Representing O2 relative to the IW buffer does not necessitate that the atmosphere of K2-18b is actually buffered by the Fe–FeO reaction at present day. Furthermore, it should be appreciated that imposing a buffered oxygen constraint does not in itself constrain the total oxygen abundance. This is evident, for example, by considering the following redox reaction:
@@
where is temperature and fugacity, in which the later can be considered equivalent to partial pressure for this discussion. If O2 is constrained relative to the IW buffer, then both O2 and are solely functions of temperature and these collectively govern the ratio of H2 to H2O but importantly not their absolute values. Imposing an additional hydrogen abundance constraint (or alternatively pressure constraint) is required to break the degeneracy to allow a unique determination of H2 and H2O. That is, two equations (Equation E3 and hydrogen mass balance) can be solved to determine two unknowns (H2 and H2O), implicitly constraining the total oxygen abundance. Hence, if O2 is prescribed as a constraint, Atmodeller solves for the total oxygen abundance self-consistently as part of the solution process. For the cooled atmosphere calculations of TRAPPIST-1e, the elemental abundance of species in the atmosphere (including oxygen) are prescribed as constraints, in which case Atmodeller then ensures a self-consistent determination of O2.
Appendix F Interplay of Solubility and Nonideality
Chachan & Stevenson (2018) and Kite et al. (2019) present models for H2 dissolution in sub-Neptunes, for ideal and nonideal scenarios, respectively. Here, we develop a toy model to demonstrate how these previous results are also obtained with the formulation of Atmodeller. We start with a simplified mass balance equation used by Atmodeller (e.g., Appendix C, Bower et al., 2022), where the first term is the dissolved mass of volatile, the second term the mass of volatile in the atmosphere, and the third term the total mass of volatile
@@ Appendix F Interplay of Solubility and Nonideality
We employ a general expression for solubility with a power-law dependence on fugacity
where , , and are constants specific to a particular equilibrium between a gas species and its dissolved component and is temperature at the atmosphere-interior interface. We define a functional form for the fugacity coefficient (Equation 1) that captures the general tendency of fugacity coefficients to increase exponentially with pressure
where is total pressure, is a constant that regulates nonideality and bar is the pressure at which the volatile behaves ideally. Hence ideal behavior is recovered as or . For simplicity, temperature dependence can also be accounted for through the choice of P0. Substituting Equations F2 and F3 into F1 gives
where
@@
The ratio of the terms in the square brackets corresponds to the ratio of the dissolved volatile mass and the atmospheric volatile mass
where is the gravitational constant. For the limit of ideal behavior () and a fully molten core () the scaling in Equation 12 of Chachan & Stevenson (2018) is recovered. Hence, the dissolved volatile mass fraction relative to that of the atmosphere increases with planet size and higher surface temperature (Figure 1, Chachan & Stevenson, 2018; Gillmann et al., 2024). Kite et al. (2019) additionally include the nonideal dissolution of H2 into magma driven by the nonlinear increase of fugacity with pressure (quantified by ), which again gives rise to enhanced volatile storage in the interior.
@@ Appendix G Additional K2-18b Figures
Figures G1 and G2 show how atmospheric pressures (a), atmospheric C/H and C/O ratios by mass (b), solubilities (c), and fugacity coefficients (d) vary with oxygen fugacity and total C/H mass ratio, respectively, between the real and ideal cases. In these simulations, the planet mass, surface radius, and temperature are the same as those in Section 3.2 and Figures 6 and 7, and only one input parameter is varied (either oxygen fugacity or C/H mass ratio), while the others are kept constant.